OpenStudy (anonymous):

Use the functions a(x) = 3x + 10 and b(x) = 2x − 8 to complete the function operations listed below. Part A: Find (a + b)(x). Show your work. (3 points) Part B: Find (a ⋅ b)(x). Show your work. (3 points) Part C: Find a[b(x)]. Show your work. (4 points)

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (anonymous):

@triciaal

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{a(x) = 3x + 10},\qquad \color{orangered}{b(x) = 2x − 8}\] \[\large\rm (a+b)(x)=\color{royalblue}{a(x)}+\color{orangered}{b(x)}\]\[\large\rm (a+b)(x)=\color{royalblue}{3x+10}+\color{orangered}{2x-8}\]Combine like-terms! :)

OpenStudy (anonymous):

5x+2?

zepdrix (zepdrix):

Yay good job! That takes care of part A.

OpenStudy (anonymous):

:}

zepdrix (zepdrix):

For part B, you'll want to put some brackets around your a(x) and b(x) so the multiplication works out properly.\[\large\rm (a\cdot b)(x)=(\color{royalblue}{a(x)})\cdot(\color{orangered}{b(x)})\]\[\large\rm (a\cdot b)(x)=(\color{royalblue}{3x+10})\cdot(\color{orangered}{2x-8})\]

OpenStudy (anonymous):

(3x+10)(2x+−8) =(3x)(2x)+(3x)(−8)+(10)(2x)+(10)(−8) =6x2−24x+20x−80 =6x2−4x−80

OpenStudy (anonymous):

Correct?

zepdrix (zepdrix):

Ooo looks good! :O

zepdrix (zepdrix):

How bout part C, the composition. Having any trouble with that one?

OpenStudy (anonymous):

Yes thats actually the one i am confused on >_<

zepdrix (zepdrix):

\[\large\rm a(\color{#DD4747}{x}) = 3\color{#DD4747}{x} + 10\]We're replacing all of the x's in our function a(x) with another function b(x).\[\large\rm a(\color{#DD4747}{b(x)}) = 3\color{#DD4747}{b(x)} + 10\]

zepdrix (zepdrix):

But again, since we have some multiplication going on, let's place brackets around the b(x) before we plug it in.

zepdrix (zepdrix):

\[\large\rm a(\color{#DD4747}{b(x)}) = 3(\color{#DD4747}{b(x)}) + 10\]\[\large\rm a(\color{#DD4747}{b(x)}) = 3(\color{#DD4747}{2x-8}) + 10\]

OpenStudy (anonymous):

=(3)(2x)+(3)(−8)+10 =6x+−24+10 =6x+−24+10 =(6x)+(−24+10) =6x+−14

zepdrix (zepdrix):

yay good job!

OpenStudy (anonymous):

i need help on part A!!!

OpenStudy (anonymous):

@enchanted_bubbles

OpenStudy (anonymous):

@zeddrix

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Then scroll up and read silly! :D

zepdrix (zepdrix):

Notice the colors

zepdrix (zepdrix):

Recall that addition is commutative, meaning we can move things around. So \(\large\rm 3x+10+2x-8\) is the same as \(\large\rm 3x+2x+10-8\) Combine your x's. You have 3 of them, and 2 more of them. 3 apples + 2 apples = ?